# Multi-modal atmospheric characterization of Pictoris b

### v1

@@ 7.1 The multi-modal analysis: advantages and limitation
7.2 On the atmospheric properties of Pictoris b
Across our study, Exo-REM proved to be the most suitable grid for modeling Pic b’s data.
@@ 7.2 On the atmospheric properties of Pictoris b
Using only GRAVITY in Sect. 4, the two preferred models, Exo-REM and Sonora, predict similar values for ( and K) and log(g) ( and dex). These results are consistent with previous atmospheric and evolutionary model predictions (see Table 12). However, Exo-REM estimates a radius that is about 10% larger than Sonora’s: compared to . By combining these radii and surface gravities, it is possible to estimate the planet’s mass. Exo-REM provides a mass estimate of , while Sonora gives ; both of which are significantly higher than the more robust dynamical measurements from Lacour et al. (2021) of . This tendency of forward models to under or overestimate both radius and surface gravity has been observed in other recent works (Palma-Bifani et al., 2023) and may stem from incorrect predictions of the emergent flux by the models after rescaling.
@@
The retrieved metallicity using only GRAVITY is super-solar for both Exo-REM () and Sonora (). A high metallicity, combined with thick cloud coverage as indicated by the very low sedimentation efficiency in Sonora (f) and previous photometric surveys (Currie et al., 2013; Bonnefoy et al., 2013; Morzinski et al., 2015; Chilcote et al., 2017; GRAVITY Collaboration et al., 2020; Stolker et al., 2020), is a surprising result, considering that clouds typically sequester a significant amount of metallic species (e.g., Mg, Si, Fe, O), which tends to lower the observed metallicity. In the case of Exo-REM, the metallicity is mainly driven by H2O absorption features while Mg, Si and Fe condense. The likely explanation for this issue (clouds + high metallicity) thus probably comes from an incorrect/imperfect reproduction of the continuum by the model. To demonstrate this, we repeated the inversion presented in Sect. 4, this time removing the continuum from the GRAVITY data. The results, shown in Table 8 below, show a noticeable change in every parameter, particularly in [M/H], which shifts from to . This lower metallicity is in fact more aligned with most planet formation model predictions, as shown on the top panel of Fig. 12. This imperfect reproduction of the continuum may stems from and incomplete and/or imperfect cloud and chemistry modeling (Rajan et al., 2017; Mollière et al., 2020): Firstly, it is possible that the collision-induced absorption of H2, which is the dominant absorber at the pressure levels probed by the K-band, is not properly modeled. Conversely, it is possible that for some spectra in the grid, the emitted flux passing through the cloud deck is not properly spectrally redistributed which can strongly modify the shape of the K-band. In both cases, the impact on the K-band shape can be sufficient enough to explain this discrepancy. Super-solar metallicities were also found by GRAVITY Collaboration et al. (2020), Worthen et al. (2024) and Houllé et al. (2024, subm.) using Exo-REM on the combined GRAVITY and available LRS data. More recently, Parker et al. (2024) detected an SiO signal in the M band using CRIRES+, further complicating the tension between clouds and metallicity, as SiO is expected to be readily sequestered into clouds (see Sect. 4.2 and Fig. 4).
| Exo-REM | (K) | log(g) (dex) | [M/H] | C/O |
@@
Conversely, Sect. 5 and 6 revealed a significantly more complex picture of Pic b’s atmosphere using a self-consistent and multi-wavelength modeling of its spectra. As previously discussed, all atmospheric parameters are extremely sensitive to the considered data, but two main regimes seem to emerge for both and log(g): a ”low-/ high-log(g)” regime ( 1500–1600 K / log(g) 4.0-5.0 dex) as predicted by Exo-REM and Sonora, and a ”high-/ low-log(g)” regime ( ¿ 1800 K / log(g) ¡ 3.5 dex) as predicted by ATMO and BT-Settl. This trend is reflected in the Bayesian evidences (see Table 6), the scaling parameters (see Fig. 9) and the residuals (see Figs. 15 and 8), indicating that ATMO and BT-Settl generally struggle more to reproduce the combined dataset shape. In the ”Physically informed” setup, Exo-REM retrieved the following values: K, log(g) dex, [M/H] and C/O . Exo-REM also remains the only model capable of retrieving a dynamically consistent mass of M . This radius-log(g) pair is in relatively good agreement with hot-start predictions from evolutionary tracks, which estimate the formation age of Pic b to be 9.6 Myr, as shown on Fig. 10. However, this simple comparison suggests a much later formation for the planet, given that its host star formed Myr ago (Mamajek & Bell, 2014) which does not align with the typical protoplanetary disk dispersal timescales (2-3 Myr, Mamajek, 2009) if we believe Pic to have been formed by core accretion. A more robust way to estimate the age of Pic b is to use our derived bolometric luminosity (log(L/L⊙) dex) alongside dynamical mass predictions (M , Lacour et al., 2021) to explore the parameter space provided by evolutionary models. Using the approach outlined in Zhang (2024) combined with the models from Spiegel & Burrows (2012), we obtained a bimodal posterior distribution for the age, with one peak near 13 Myr and a secondary peak around 84 Myr (see Fig. 11). While the first solution is physically plausible, it is important to acknowledge that both the initial entropy and age remain poorly constrained. This behavior, also noted in Zhang (2024), mainly stems from the large dynamical mass uncertainties. Detailed orbital monitoring is necessary if we want to push for independent planet age estimations. A more complete discussion on formation scenarios can be found in the next section.
@@ 7.3 Interpreting formation tracers
Exo-REM and Sonora both predicts super-solar metallicities regardless of the dataset added, apart when using GRAVITY data alone with Sonora (Sect. 5). The challenges mentioned above regarding the reproduction of the continuum still holds here. We can add that, for the Y to K bands, the treatment of collisionally broadened K I doublet is also essential. In particular, Phillips et al. (2020) showed that a simple broadening model has a strong impact on the Y, J, H (SPHERE, GPI), and K-band (GRAVITY) shapes. The unprecedented wavelength coverage (0.9-7 µm) provided by all these observations also enable us to retrieve robust measurements of the bolometric luminosity (log(L/L⊙) dex, Exo-REM) which can be used to constrain the maximum metallic content of the atmosphere using simple interior models. We used the methodology presented in Marleau & Cumming (2014); Marleau et al. (2019) to compare cooling models constraints on planet’s mass and post-formation entropy with the results of population synthesis from Emsenhuber et al. (2021a, b). Assuming the metallic core is completely diluted (i.e. metallic content = atmospheric metallicity), we get an optimistic upper limit on [M/H] of 0.05. This means that the super-solar metallicity we retrieve () is hard to reach with this setup. Interaction with the disk, as discussed in the next section (top panel of Fig. 12) might help reaching a super-solar value.
We obtain a solar value for the C/O using Exo-REM, which remains robust with the addition of datasets, yielding a final value of C/O . The solar value retrieve by Exo-REM is in tension with GRAVITY Collaboration et al. (2020) (), Landman et al. (2024) () and Worthen et al. (2024) () but in good agreement with more recent studies from Kiefer et al. (2024) () and Houllé et al. (2024, subm.) (). All of these values (except those from Landman et al., 2024) were also obtained using Exo-REM, ruling out the simpler cloud prescription in the retrieval as a possible explanation for the observed discrepancies.
@@
Kiefer et al. (2024) report C/O values similar to ours and their conclusions on the formation history of Pic b therefore still hold based on simple assumptions on the disk composition. We propose here in addition to use planet population models assuming the pebble accretion scenario and encompassing the core+enveloppe formation of the planets as well as drift and evaporations from pebbles (Schneider & Bitsch, 2021a, b) to further interpret our results. Figure 12 compares planets formed with this model that fall within the same mass and semi-major axis range as Pic b with the different Exo-REM predictions presented in this work. While this model could potentially explain the higher metallicity we observe (top panel), the situation is reversed for the C/O ratio. Both the full model (Evaporating Pebbles) and the reduced model (Plain) predict C/O values greater than 1. We can notice as well that our C/O and M/H estimates might be biased because atmospheric models rarely explore values beyond this threshold. This is shown by the blue and red vertical lines in the bottom panel, representing the upper limits of Exo-REM and BT-Settl respectively.
@@ 8 Conclusion
In this work, we presented the first high-resolution spectro-interferometric observation of Pic b using VLTI/GRAVITY at R 4 000. This new dataset brings insights onto the atmosphere of this object :
The continuum in the K band is primarily shaped by H2O absorption, allowing us to place strong constraints on the effective temperature and surface gravity of the object. Although statistical metrics rank Sonora as the best overall grid, half of its intrinsic parameters reach the grid’s edges. Therefore, we adopt the second-best statistically preferred grid, Exo-REM, as our best-fit model. It yields K and log(g). Collision-induced absorption of H2 which is sensitive to atmospheric pressure, also introduce metallicity-related information into this continuum. However, discrepancies between the retrieved metallicities from this spectrum and its continuum-subtracted counterpart suggest an issue in the modeling; whether due to inaccuracies in the collision-induced absorption, deviant spectra in the grids, or other systematics which prevents us from placing robust constraints on [M/H].
We found a near-solar C/O value of , but we were only able to place a loose upper limit on log(12C/13C), likely due to the combined effects of telluric residuals, wavelength calibration issues, and/or limitations in grid sampling, as the 13C molecular features are very narrow.
We also investigated the combination of previously obtained spectro-photometric observations using our newly developed framework, MOSAIC. This framework allowed us to simultaneously invert multiple observations and explore their impact on the retrieved parameters when adding them iteratively. We found that all atmospheric parameters are significantly influenced by the dataset used, and that the models typically separate into two categories: ”Low-/ high-log(g)” ( 1500–1600 K / log(g) 4.0-5.0 dex) as predicted by Exo-REM and Sonora, and a ”high-/ low-log(g)” ( ¿ 1800 K / log(g) ¡ 3.5 dex) as predicted by ATMO and BT-Settl. The first category appears to be statistically favored and more aligned with previous studies on the target.
Adding high-resolution spectroscopy from CRIRES+ in this framework proved to be challenging (forward models resolution, inversion time, …). The shorter coverage of both the K- and M-bands observation as well as their noise limits their impact on the final fitted parameters. Finally, for the future, new data from the improved GRAVITY+ instrument could better mitigate telluric contamination in the K-band and enhance the retrieval of the 13CO signal. Meanwhile, Pic b will be a compelling target for short-term monitoring with JWST, allowing us to explore potential variability and gain new insights into its cloud coverage and properties. Additionally, the advanced capabilities of the upcoming Extremely Large Telescope (ELT) by 2030 will enable more detailed analysis including the use of Doppler imaging to map temperature and chemical inhomogeneities and clouds.
@@ Acknowledgements.
This work is heavily supported by a large collaboration of international people: Atmospheric modeling groups represented by B. Charnay (Exo-REM), P. Tremblin (ATMO), D. Homeier, F. Allard (deceased) (BT-Settl), J. Fortney, C. Morley (Sonora). Collaborators who provided us with the data are represented by the GRAVITY collab (GRAVITY high-res.), A. Chomez (SPHERE), F. Kiefer (SINFONI), M. Houllé (MATISSE), R. Landman, L T. Parker (CRIRES+) and K. Worthen (MIRI). This work is based on observations collected at the European Southern Observatory under ESO Large programme ExoGRAVITY: ID 1104.C-0651. S. Lacour acknowledges the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-21-CE31-0017 (project ExoVLTI) and grant ANR-23-CE31-0006-01 (MIRAGES). G.-D. Marleau acknowledges support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28 (PI: Th. Henning).References
- Ackerman & Marley (2001) Ackerman, A. S. & Marley, M. S. 2001, The Astrophysical Journal, 556, 872, publisher: IOP ADS Bibcode: 2001ApJ…556..872A
- Allard et al. (2012) Allard, F., Homeier, D., & Freytag, B. 2012, Philosophical Transactions of the Royal Society of London Series A, 370, 2765, aDS Bibcode: 2012RSPTA.370.2765A
- Baudino et al. (2015) Baudino, J. L., Bézard, B., Boccaletti, A., et al. 2015, Astronomy and Astrophysics, 582, A83, aDS Bibcode: 2015A&A…582A..83B
- Benneke & Seager (2013) Benneke, B. & Seager, S. 2013, The Astrophysical Journal, 778, 153, publisher: IOP ADS Bibcode: 2013ApJ…778..153B
@@
- Bonnefoy et al. (2013) Bonnefoy, M., Boccaletti, A., Lagrange, A. M., et al. 2013, Astronomy and Astrophysics, 555, A107, aDS Bibcode: 2013A&A…555A.107B
- Bonnefoy et al. (2011) Bonnefoy, M., Lagrange, A. M., Boccaletti, A., et al. 2011, Astronomy and Astrophysics, 528, L15, aDS Bibcode: 2011A&A…528L..15B
- Bonnefoy et al. (2014) Bonnefoy, M., Marleau, G. D., Galicher, R., et al. 2014, Astronomy and Astrophysics, 567, L9, aDS Bibcode: 2014A&A…567L…9B
@@
- Brandt et al. (2021) Brandt, G. M., Brandt, T. D., Dupuy, T. J., Li, Y., & Michalik, D. 2021, The Astronomical Journal, 161, 179, publisher: IOP ADS Bibcode: 2021AJ….161..179B
- Brogi & Line (2019) Brogi, M. & Line, M. R. 2019, The Astronomical Journal, 157, 114, publisher: IOP ADS Bibcode: 2019AJ….157..114B
- Charnay et al. (2018) Charnay, B., Bézard, B., Baudino, J. L., et al. 2018, The Astrophysical Journal, 854, 172, publisher: IOP ADS Bibcode: 2018ApJ…854..172C
@@
- Chilcote et al. (2015) Chilcote, J., Barman, T., Fitzgerald, M. P., et al. 2015, The Astrophysical Journal, 798, L3, publisher: IOP ADS Bibcode: 2015ApJ…798L…3C
- Chilcote et al. (2017) Chilcote, J., Pueyo, L., De Rosa, R. J., et al. 2017, The Astronomical Journal, 153, 182, publisher: IOP ADS Bibcode: 2017AJ….153..182C
@@
- Currie et al. (2013) Currie, T., Burrows, A., Madhusudhan, N., et al. 2013, The Astrophysical Journal, 776, 15, publisher: IOP ADS Bibcode: 2013ApJ…776…15C
- Cushing et al. (2008) Cushing, M. C., Marley, M. S., Saumon, D., et al. 2008, The Astrophysical Journal, 678, 1372, publisher: IOP ADS Bibcode: 2008ApJ…678.1372C
@@
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, Astronomy and Astrophysics, 674, A1, aDS Bibcode: 2023A&A…674A…1G
@@
- Gontcharov (2006) Gontcharov, G. A. 2006, Astronomy Letters, 32, 759, publisher: Springer ADS Bibcode: 2006AstL…32..759G
@@
- GRAVITY Collaboration et al. (2020) GRAVITY Collaboration, Nowak, M., Lacour, S., et al. 2020, Astronomy and Astrophysics, 633, A110, aDS Bibcode: 2020A&A…633A.110G
@@
- Hauschildt et al. (1999) Hauschildt, P. H., Allard, F., & Baron, E. 1999, The Astrophysical Journal, 512, 377, publisher: IOP ADS Bibcode: 1999ApJ…512..377H
- Hayoz et al. (2023) Hayoz, J., Cugno, G., Quanz, S. P., et al. 2023, Astronomy and Astrophysics, 678, A178, aDS Bibcode: 2023A&A…678A.178H
@@
- Hoeijmakers et al. (2018) Hoeijmakers, H. J., Schwarz, H., Snellen, I. A. G., et al. 2018, Astronomy and Astrophysics, 617, A144, aDS Bibcode: 2018A&A…617A.144H
@@
- Kiefer et al. (2024) Kiefer, F., Bonnefoy, M., Charnay, B., et al. 2024, Astronomy and Astrophysics, 685, A120, aDS Bibcode: 2024A&A…685A.120K
@@
- Knuth et al. (2015) Knuth, K. H., Habeck, M., Malakar, N. K., Mubeen, A. M., & Placek, B. 2015, Digital Signal Processing, 47, 50, aDS Bibcode: 2015DSP….47…50K
@@
- Lacour et al. (2021) Lacour, S., Wang, J. J., Rodet, L., et al. 2021, Astronomy and Astrophysics, 654, L2, aDS Bibcode: 2021A&A…654L…2L
- Lagrange et al. (2009) Lagrange, A. M., Gratadour, D., Chauvin, G., et al. 2009, Astronomy and Astrophysics, 493, L21, aDS Bibcode: 2009A&A…493L..21L
- Lagrange et al. (2020) Lagrange, A. M., Rubini, P., Nowak, M., et al. 2020, Astronomy and Astrophysics, 642, A18, aDS Bibcode: 2020A&A…642A..18L
- Landman et al. (2024) Landman, R., Stolker, T., Snellen, I. A. G., et al. 2024, Astronomy and Astrophysics, 682, A48, aDS Bibcode: 2024A&A…682A..48L
@@
- Males et al. (2014) Males, J. R., Close, L. M., Morzinski, K. M., et al. 2014, The Astrophysical Journal, 786, 32, publisher: IOP ADS Bibcode: 2014ApJ…786…32M
@@
- Mamajek & Bell (2014) Mamajek, E. E. & Bell, C. P. M. 2014, Monthly Notices of the Royal Astronomical Society, 445, 2169, publisher: OUP ADS Bibcode: 2014MNRAS.445.2169M
@@
- Millar-Blanchaer et al. (2015) Millar-Blanchaer, M. A., Graham, J. R., Pueyo, L., et al. 2015, The Astrophysical Journal, 811, 18, publisher: IOP ADS Bibcode: 2015ApJ…811…18M
@@
- Morris et al. (2024) Morris, E. C., Wang, J. J., Hsu, C.-C., et al. 2024, arXiv e-prints, arXiv:2405.13125
- Morzinski et al. (2015) Morzinski, K. M., Males, J. R., Skemer, A. J., et al. 2015, The Astrophysical Journal, 815, 108, publisher: IOP ADS Bibcode: 2015ApJ…815..108M
@@
- Nasedkin et al. (2024) Nasedkin, E., Mollière, P., Lacour, S., et al. 2024, arXiv e-prints, arXiv:2404.03776
- Nissen (2013) Nissen, P. E. 2013, Astronomy and Astrophysics, 552, A73, aDS Bibcode: 2013A&A…552A..73N

### v2

@@ 7.1 The multimodal analysis: Advantages and limitations
7.2 Exploring the atmospheric properties of Pictoris b
@@
Across our study, Exo-REM proved to be the most suitable grid for modeling Pic b’s data. Using only GRAVITY in Sect. 4, the two preferred models, Exo-REM and Sonora, predict similar values for ( and K) and log(g) ( and dex). These results are consistent with previous atmospheric and evolutionary model predictions (see Table 15). However, Exo-REM estimates a radius that is about 10% larger than Sonora’s: compared to . By combining these radii and surface gravities, it is possible to estimate the planet’s mass. Exo-REM provides a mass estimate of , while Sonora gives ; both of which are significantly higher than the more robust dynamical measurements from Lacour et al. (2021) of . This tendency of forward models to under or overestimate both radius and surface gravity has been observed in other recent works (Palma-Bifani et al., 2023) and could stem from incorrect predictions of the emergent flux by the models after rescaling.
@@
The retrieved metallicity using only GRAVITY is super-solar for both Exo-REM () and Sonora (). A high metallicity, combined with thick cloud coverage as indicated by the very low sedimentation efficiency in Sonora (f) and previous photometric surveys (Currie et al., 2013; Bonnefoy et al., 2013; Morzinski et al., 2015; Chilcote et al., 2017; GRAVITY Collaboration et al., 2020; Stolker et al., 2020), is a surprising result, considering that clouds typically sequester a significant amount of metallic species (e.g., Mg, Si, Fe, O), which tends to lower the observed metallicity. In the case of Exo-REM, the metallicity is mainly driven by H2O absorption features while Mg, Si and Fe condense. The likely explanation for this issue (clouds + high metallicity) is likely to come from an incorrect or imperfect reproduction of the continuum by the model. To demonstrate this, we repeated the inversion presented in Sect. 4, this time removing the continuum from the GRAVITY data. The results, shown in Table 8 below, show a noticeable change in every parameter, particularly in [M/H], which shifts from to . This lower metallicity is in fact more aligned with most planet formation model predictions, as shown on the top panel of Fig. 12. This imperfect reproduction of the continuum may stems from and incomplete and/or imperfect cloud and chemistry modeling (Rajan et al., 2017; Mollière et al., 2020): Firstly, it is possible that the collision-induced absorption of H2, which is the dominant absorber at the pressure levels probed by the K band, is not properly modeled. Conversely, it is possible that for some spectra in the grid, the emitted flux passing through the cloud deck is not properly spectrally redistributed which can strongly modify the shape of the K band. In both cases, the impact on the K-band shape can be sufficient enough to explain this discrepancy. Super-solar metallicities were also found by GRAVITY Collaboration et al. (2020), Worthen et al. (2024) and Houllé et al. (2025) using Exo-REM on the combined GRAVITY and available LRS data. More recently, Parker et al. (2024) detected an SiO signal in the M band using CRIRES+, further complicating the tension between clouds and metallicity, as SiO is expected to be readily sequestered into clouds (see Sect. 4.2 and Fig. 4).
| Exo-REM | log(g) | [M/H] | C/O | |
| (K) | (dex) | |||
@@
Conversely, Sect. 5 and 6 revealed a significantly more complex picture of Pic b’s atmosphere using a self-consistent and multi-wavelength modeling of its spectra. As previously discussed, all atmospheric parameters are extremely sensitive to the considered data, but two main regimes seem to emerge for both and log(g): a ”low-/ high-log(g)” regime ( 1500–1600 K / log(g) 4.0–5.0 dex) as predicted by Exo-REM and Sonora, and a ”high-/ low-log(g)” regime ( ¿ 1800 K / log(g) ¡ 3.5 dex) as predicted by ATMO and BT-Settl. This trend is reflected in the Bayesian evidences (see Table 12), the scaling parameters (see Fig. 9) and the residuals (see Figs. 15 and 8), indicating that ATMO and BT-Settl generally struggle more to reproduce the combined dataset shape. In the ”physically informed” setup, Exo-REM retrieved the following values: K, log(g) dex, [M/H] and C/O . Exo-REM also remains the only model capable of retrieving a dynamically consistent mass of M . This radius-log(g) pair is in relatively good agreement with hot-start predictions from evolutionary tracks, which estimate the formation age of Pic b to be 9.6 Myr, as shown on Fig. 10. However, this simple comparison suggests a much later formation for the planet, given that its host star formed Myr ago (Mamajek & Bell, 2014) which does not align with the typical protoplanetary disk dispersal timescales (2–3 Myr, Mamajek, 2009) if we believe Pic to have been formed by core accretion. A more robust way to estimate the age of Pic b is to use our derived bolometric luminosity (log(L/L⊙) dex) alongside dynamical mass predictions (M , Lacour et al., 2021) to explore the parameter space provided by evolutionary models. Using the approach outlined in Zhang (2024) combined with the models from Spiegel & Burrows (2012), we obtained a bimodal posterior distribution for the age, with one peak near 13 Myr and a secondary peak around 84 Myr (see Fig. 11). While the first solution is physically plausible, it is important to acknowledge that both the initial entropy and age remain poorly constrained. This behavior, also noted in Zhang (2024), mainly stems from the large dynamical mass uncertainties. Detailed orbital monitoring is necessary if we want to push for independent planet age estimations. A more complete discussion on formation scenarios can be found in the next section.
@@ 7.3 Interpreting formation tracers
Exo-REM and Sonora both predicts super-solar metallicities regardless of the dataset added, apart when using GRAVITY data alone with Sonora (Sect. 5). The challenges mentioned above regarding the reproduction of the continuum still holds here. We can add that, for the Y to K bands, the treatment of collisionally broadened K I doublet is also essential. In particular, Phillips et al. (2020) showed that a simple broadening model has a strong impact on the Y-, J-, H- (SPHERE, GPI), and K-band (GRAVITY) shapes. The unprecedented wavelength coverage (0.9–7 µm) provided by all these observations also enable us to retrieve robust measurements of the bolometric luminosity (log(L/L⊙) dex, Exo-REM), which could be used to constrain the maximum metallic content of the atmosphere using simple interior models. We used the methodology presented in Marleau & Cumming (2014); Marleau et al. (2019) to compare cooling models constraints on planet’s mass and post-formation entropy with the results of population synthesis from Emsenhuber et al. (2021a, b). Assuming the metallic core is completely diluted (i.e., metallic content = atmospheric metallicity), we get an optimistic upper limit on [M/H] of 0.05. This means that the super-solar metallicity we retrieve () is hard to reach with this setup. Interaction with the disk, as discussed in the next section (top panel of Fig. 12) might help reaching a super-solar value.
@@
We obtain a solar value for the C/O using Exo-REM, which remains robust with the addition of datasets, yielding a final value of C/O . The solar value retrieve by Exo-REM is in tension with GRAVITY Collaboration et al. (2020) (), Landman et al. (2024) () and Worthen et al. (2024) () but in good agreement with more recent studies from Kiefer et al. (2024) () and Houllé et al. (2025) (). All of these values (except those from Landman et al., 2024) were also obtained using Exo-REM, ruling out the simpler cloud prescription in the retrieval as a possible explanation for the observed discrepancies. Kiefer et al. (2024) report C/O values similar to ours and their conclusions on the formation history of Pic b therefore still hold based on simple assumptions on the disk composition. We propose here the additional use of planet population models, assuming the pebble accretion scenario and encompassing the core+envelope formation of the planets as well as drift and evaporations from pebbles (Schneider & Bitsch, 2021a, b) to further interpret our results. Figure 12 compares planets formed with this model that fall within the same mass and semi-major axis range as Pic b with the different Exo-REM predictions presented in this work. While this model could potentially explain the higher metallicity we observe (top panel), the situation is reversed for the C/O ratio. Both the full model (evaporating pebbles) and the reduced model (plain) predict C/O values greater than 1. We can notice as well that our C/O and M/H estimates might be biased because atmospheric models rarely explore values beyond this threshold. This is shown by the blue and red vertical lines in the bottom panel, representing the upper limits of Exo-REM and BT-Settl respectively.
@@ 8 Conclusion
In this work, we present the first high-resolution spectro-interferometric observation of Pic b using VLTI/GRAVITY at R 4,000. This new dataset brings insights onto the atmosphere of this object. First, the continuum in the K band is primarily shaped by H2O absorption, allowing us to place strong constraints on the effective temperature and surface gravity of the object. Although statistical metrics rank Sonora as the best overall grid, half of its intrinsic parameters reach the grid’s edges. Therefore, we adopted the second-best statistically preferred grid, Exo-REM, as our best-fit model. It yields K and log(g). Collision-induced absorption of H2 which is sensitive to atmospheric pressure, also introduce metallicity-related information into this continuum. However, discrepancies between the retrieved metallicities from this spectrum and its continuum-subtracted counterpart suggest an issue in the modeling; either due to inaccuracies in the collision-induced absorption, deviant spectra in the grids, or other systematics that prevents us from placing robust constraints on [M/H].
We found a near-solar C/O value of , but we were only able to place a loose upper limit on log(12C/13C), likely due to the combined effects of telluric residuals, wavelength calibration issues, and/or limitations in grid sampling, as the 13C molecular features are very narrow. We also investigated the combination of previously obtained spectro-photometric observations using our newly developed framework, MOSAIC. This framework allowed us to simultaneously invert multiple observations and explore their impact on the retrieved parameters when adding them iteratively. We found that all atmospheric parameters are significantly influenced by the dataset used, and that the models typically separate into two categories: ”low-/ high-log(g)” ( 1500–1600 K / log(g) 4.0–5.0 dex) as predicted by Exo-REM and Sonora, and a ”high-/ low-log(g)” ( ¿ 1800 K / log(g) ¡ 3.5 dex) as predicted by ATMO and BT-Settl. The first category appears to be statistically favored and more aligned with previous studies on the target.
Adding high-resolution spectroscopy from CRIRES+ in this framework proved to be challenging (forward models resolution, inversion time, etc.). The shorter coverage of both the K- and M-band observations, as well as their noise, limits their impact on the final fitted parameters. Finally, for the future, new data from the improved GRAVITY+ instrument could better mitigate telluric contamination in the K band and enhance the retrieval of the 13CO signal. Meanwhile, Pic b will be a compelling target for short-term monitoring with JWST, allowing us to explore potential variability and gain new insights into its cloud coverage and properties. Additionally, the advanced capabilities of the upcoming Extremely Large Telescope (ELT) by 2030 will enable more detailed analyses, including the use of Doppler imaging to map temperature and chemical inhomogeneities and clouds.
@@ Acknowledgements.
This work is heavily supported by a large collaboration of international people: Atmospheric modeling groups represented by B. Charnay (Exo-REM), P. Tremblin (ATMO), D. Homeier, F. Allard (deceased) (BT-Settl), J. Fortney, C. Morley (Sonora). Collaborators who provided us with the data are represented by the GRAVITY collab (GRAVITY high-res.), A. Chomez (SPHERE), F. Kiefer (SINFONI), M. Houllé (MATISSE), R. Landman, L T. Parker (CRIRES+) and K. Worthen (MIRI). This work is based on observations collected at the European Southern Observatory under ESO Large programme ExoGRAVITY: ID 1104.C-0651. S. Lacour acknowledges the support of the French Agence Nationale de la Recherche (ANR), under grant ANR-21-CE31-0017 (project ExoVLTI) and grant ANR-23-CE31-0006-01 (MIRAGES). G.-D. Marleau acknowledges support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28 (PI: Th. Henning). This work is based [in part] on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with GTO program 1294.References
- Ackerman & Marley (2001) Ackerman, A. S. & Marley, M. S. 2001, ApJ, 556, 872
- Allard et al. (2012) Allard, F., Homeier, D., & Freytag, B. 2012, Philosophical Transactions of the Royal Society of London Series A, 370, 2765
- Baudino et al. (2015) Baudino, J. L., Bézard, B., Boccaletti, A., et al. 2015, A&A, 582, A83
- Benneke & Seager (2013) Benneke, B. & Seager, S. 2013, ApJ, 778, 153
@@
- Bonnefoy et al. (2013) Bonnefoy, M., Boccaletti, A., Lagrange, A. M., et al. 2013, A&A, 555, A107
- Bonnefoy et al. (2011) Bonnefoy, M., Lagrange, A. M., Boccaletti, A., et al. 2011, A&A, 528, L15
- Bonnefoy et al. (2014) Bonnefoy, M., Marleau, G. D., Galicher, R., et al. 2014, A&A, 567, L9
@@
- Brandt et al. (2021) Brandt, G. M., Brandt, T. D., Dupuy, T. J., Li, Y., & Michalik, D. 2021, AJ, 161, 179
- Brogi & Line (2019) Brogi, M. & Line, M. R. 2019, AJ, 157, 114
- Charnay et al. (2018) Charnay, B., Bézard, B., Baudino, J. L., et al. 2018, ApJ, 854, 172
@@
- Chilcote et al. (2015) Chilcote, J., Barman, T., Fitzgerald, M. P., et al. 2015, ApJ, 798, L3
- Chilcote et al. (2017) Chilcote, J., Pueyo, L., De Rosa, R. J., et al. 2017, AJ, 153, 182
@@
- Currie et al. (2013) Currie, T., Burrows, A., Madhusudhan, N., et al. 2013, ApJ, 776, 15
- Cushing et al. (2008) Cushing, M. C., Marley, M. S., Saumon, D., et al. 2008, ApJ, 678, 1372
@@
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
@@
- Gontcharov (2006) Gontcharov, G. A. 2006, Astronomy Letters, 32, 759
@@
- GRAVITY Collaboration et al. (2020) GRAVITY Collaboration, Nowak, M., Lacour, S., et al. 2020, A&A, 633, A110
@@
- Hauschildt et al. (1999) Hauschildt, P. H., Allard, F., & Baron, E. 1999, ApJ, 512, 377
- Hayoz et al. (2023) Hayoz, J., Cugno, G., Quanz, S. P., et al. 2023, A&A, 678, A178
@@
- Hoeijmakers et al. (2018) Hoeijmakers, H. J., Schwarz, H., Snellen, I. A. G., et al. 2018, A&A, 617, A144
- Houllé et al. (2025) Houllé, M., Millour, F., Berio, P., et al. 2025, arXiv e-prints, arXiv:2508.18366
@@
- Kiefer et al. (2024) Kiefer, F., Bonnefoy, M., Charnay, B., et al. 2024, A&A, 685, A120
@@
- Knuth et al. (2015) Knuth, K. H., Habeck, M., Malakar, N. K., Mubeen, A. M., & Placek, B. 2015, Digital Signal Processing, 47, 50
@@
- Lacour et al. (2021) Lacour, S., Wang, J. J., Rodet, L., et al. 2021, A&A, 654, L2
- Lagrange et al. (2009) Lagrange, A. M., Gratadour, D., Chauvin, G., et al. 2009, A&A, 493, L21
- Lagrange et al. (2020) Lagrange, A. M., Rubini, P., Nowak, M., et al. 2020, A&A, 642, A18
- Landman et al. (2024) Landman, R., Stolker, T., Snellen, I. A. G., et al. 2024, A&A, 682, A48
@@
- Males et al. (2014) Males, J. R., Close, L. M., Morzinski, K. M., et al. 2014, ApJ, 786, 32
@@
- Mamajek & Bell (2014) Mamajek, E. E. & Bell, C. P. M. 2014, MNRAS, 445, 2169
@@
- Millar-Blanchaer et al. (2015) Millar-Blanchaer, M. A., Graham, J. R., Pueyo, L., et al. 2015, ApJ, 811, 18
@@
- Morris et al. (2024) Morris, E. C., Wang, J. J., Hsu, C.-C., et al. 2024, AJ, 168, 144
- Morzinski et al. (2015) Morzinski, K. M., Males, J. R., Skemer, A. J., et al. 2015, ApJ, 815, 108
@@
- Nasedkin et al. (2024) Nasedkin, E., Mollière, P., Lacour, S., et al. 2024, A&A, 687, A298
- Nissen (2013) Nissen, P. E. 2013, A&A, 552, A73